5. The polynomial equation x (x + 1) + 8 = (x + 2) {x – 2) is
(a) linear equation
(b) quadratic equation
(c) cubic equation
(d) bi-quadratic equation

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: a Explaination:Reason: We have x(x + 1) + 8 = (x + 2) (x – 2) ⇒ x² + x + 8 = x² – 4 ⇒ x² + x + 8- x² + 4 = 0 ⇒ x + 12 = 0, which is a linear equation.